Two former Northwest University students work in investment banks with salaries of 560,000 each for 2 years after they graduate. together they saved s50,000. after 2 years, they decided to quit their job and start a website designing business. they used the 50,000 to buy computer equipment, tables and chairs. for the next 2 years, they earn s40,000 annually, pay themselves s10,000 annually, and rent an office for s18,000 per year,. Prior to the investment, their S50,000 was in bonds with 10 percent interest. are they now getting economic benefits? explain your answer




1
Expert's answer
2021-10-25T08:47:39-0400

Given,

Salary of the students in the investment bank "= 560,000"

Graduation has been done after 2 years

Saved amount after 2 year "= 50000"

Annual earning after 2 years "= 40,000"

Pay to themselves "= 10000"

Office rent"= 18000" per year

Amount which is in bond "= 50000"

Interest on bond "= 10\%"

We know that, economic profit is kind of profit in which the excess of the opportunity cost of the money they spend to produce their product.

The opportunity cost of their time each per year "= 560000\/2 = 280000"

Opportunity cost of the capital on bod = "50000\times \frac{10}{100} =5000"

Hence, total opportunity cost "= 560000+5000 =565000"

Net profit "= 40000-18000 =22000"

Economic profit = Net profit - Opportunity cost

"= 22000-565000 = 543000"

Net economic loss "= 545000"